Offering 1,500 points for each problem. And will add additional points if everything is worked out and correct. Thank you, and have a blessed day.
Answer #1

Malus Law:

I = I0 * (cos60)^2 = I0 / 4
Answer #2

by maulas law

I = Iocos^2(thetha)

I = Iocos^2(30)

I = 0.75Io

75%
Answer #3

I*cos^2(teta)

I through 1

I/4 through 2

then 25% passes through